# Long-Range Pairing in the Kitaev Model: Krylov Subspace Signatures

### v2

@@ B.4 Reorthogonalization and stability checks
Figure 5: BdG spectrum $E_{\nu}/\sqrt{\mathrm{Tr}(H^{2})}$ versus $\theta/\pi$ at extreme pairing-dominated $\epsilon=10$ for three long-range exponents $\alpha$ ($N=1000$, open boundaries). The systematic $\alpha$-dependence observed in Figs. 1 and 4 persists: gapless region for $\alpha=3$, similar gapless range for $\alpha=1$ with gap opening, and degeneracy lifted except for a small near-degenerate regime for $\alpha=1/3$. Spectral density near $E=0$ is further increased relative to smaller $\epsilon$, occurring uniformly across all $\alpha$ including the short-range limit. The qualitative $\alpha$-dependence remains unchanged across various $\epsilon$ regimes.
Appendix C Additional Results for Spectrum and the Role of Long-Range Pairing
@@
Figure 4 displays the spectrum at $\epsilon=1$. The qualitative $\alpha$-dependence observed at $\epsilon=-0.2$ (Fig. 1) persists: panel 4 ($\alpha=3$) exhibits a gapless region centered near $\theta/\pi\approx 0.5$ with sparse spectral density near $E=0$ elsewhere; panel 4 ($\alpha=1$) shows modified spectral density; and panel LABEL:fig:spectrum_alpha01_eps_1 ($\alpha=0.1$) displays the degeneracy lifted except for a small near-degenerate regime, with markedly denser spectral density near $E=0$ throughout. Compared to $\epsilon=-0.2$, the enhanced pairing amplitude at $\epsilon=1$ increases the density of modes near $E=0$ across all three $\alpha$ values.
Figure 5 presents the extreme pairing-dominated limit $\epsilon=10$. The systematic $\alpha$-dependence remains qualitatively unchanged: gapless region over a finite $\theta/\pi$ range for $\alpha=3$ (panel 5), similar gapless range with gap opening for $\alpha=1$ (panel 5), and degeneracy lifting to a small near-degenerate regime for $\alpha=1/3$ (panel LABEL:fig:spectrum_alpha01_eps_10). The spectral density near $E=0$ is further amplified relative to both $\epsilon=-0.2$ and $\epsilon=1$, reflecting the dominance of the pairing term $(1+\epsilon)\sin\theta$ in Eq. (7b). Notably, this increased density occurs uniformly across all $\alpha$ regimes, including the short-range limit where exponential localization would otherwise suppress accumulation of states near $E=0$.
@@
Figure 6: Lanczos coefficients $\{b_{n}\}$ for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries and boundary seed $\gamma_{1}+\gamma_{2}$ ($N=1000$). Each panel shows the two interleaved subsequences (odd and even recursion steps), whose relative ordering determines the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ and crossing count $N_{\mathrm{cross}}$. Panels 6 ($\alpha=2$, $\theta/\pi=0.1$) and 6 ($\alpha=2/3$, $\theta/\pi=0.1$) lie in the bulk-gap regime and exhibit no interchange of the two subsequences ($N_{\mathrm{cross}}=0$), while panels 6 ($\alpha=2$, $\theta/\pi=0.4$) and 6 ($\alpha=2/3$, $\theta/\pi=0.4$) lie in the edge-gap regime and show clear interchanges ($N_{\mathrm{cross}}\geq 1$). The qualitative pattern of no crossing in the bulk-gap phase as well as nonzero crossings in the edge-gap phase matches that observed for the single-operator seed $\gamma_{1}$ (Fig. 2), though the quantitative sharpness of the phase diagram when scanned over full parameter space $(\alpha,\theta)$ is reduced compared to the minimal boundary seed presented in the main text (compare Fig. 3 and Fig. 9). As in Fig. 2, the total number of Lanczos coefficients varies across parameter points. To ensure numerical stability, the recursion is terminated when $b_{n}\lesssim 10^{-7}$, and all subsequent coefficients are excluded from the analysis (see Appendix B.4 for details on stability checks).
Appendix D Alternative seed operators
@@
Figures 6, 7, and 8 show the Lanczos coefficients for these three seeds at representative parameter points spanning both short-range and long-range regimes, as well as bulk-gap and edge-gap phases. In all cases, the qualitative pattern observed in the main text persists: the two interleaved subsequences (odd and even recursion steps) do not interchange in the bulk-gap regime, yielding $N_{\mathrm{cross}}=0$, while clear interchanges occur in the edge-gap regime, producing $N_{\mathrm{cross}}\geq 1$. This confirms that the crossing-count signature is robust across seed choices and reflects genuine edge-bulk physics rather than an artifact of the specific seed operator.
However, when the crossing-count diagnostic is computed over the full parameter space $(\alpha,\theta)$, the quantitative sharpness of the phase boundary varies systematically with seed localization, as shown in Fig. 9. The boundary seed $\gamma_{1}+\gamma_{2}$ (panel 9) produces a phase diagram in close quantitative agreement with the BdG-derived edge-bulk gap boundary, though slightly less sharp than the minimal seed $\gamma_{1}$ used in the main text. By contrast, the bulk seeds $\gamma_{N}$ and $\gamma_{N}+\gamma_{N+1}$ (panels 9, 9) couple comparably to both chain ends and predominantly probe bulk-extended excitations, resulting in noticeably degraded quantitative matching with the BdG boundary even though the qualitative edge-bulk distinction remains visible. These results demonstrate that boundary-localized seeds are essential for obtaining a quantitatively reliable Krylov diagnostic of edge-versus-bulk gap control, while bulk seeds provide at best a qualitative probe.
@@
Figure 7: Lanczos coefficients $\{b_{n}\}$ for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries and bulk seed $\gamma_{N}$ ($N=1000$). Each panel shows the two interleaved subsequences (odd and even recursion steps), whose relative ordering determines the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ and crossing count $N_{\mathrm{cross}}$. Panels 7 ($\alpha=2$, $\theta/\pi=0.1$) and 7 ($\alpha=2/3$, $\theta/\pi=0.1$) lie in the bulk-gap regime ($N_{\mathrm{cross}}=0$), while panels 7 ($\alpha=2$, $\theta/\pi=0.4$) and 7 ($\alpha=2/3$, $\theta/\pi=0.4$) lie in the edge-gap regime ($N_{\mathrm{cross}}\geq 1$). The qualitative edge-bulk distinction persists for this bulk-localized seed, which couples comparably to both chain ends, though the phase diagram (obtained by scanning the full parameter space $(\alpha,\theta)$) sharpness degrades compared to boundary seeds (compare Fig. 3 and Fig. 9). To maintain numerical rigor, the recursion is terminated when $b_{n}\lesssim 10^{-7}$, excluding all subsequent coefficients (see Appendix B.4).
@@
Figure 8: Lanczos coefficients $\{b_{n}\}$ for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries and bulk seed $\gamma_{N}+\gamma_{N+1}$ ($N=1000$). Each panel shows the two interleaved subsequences whose relative ordering determines the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ and crossing count $N_{\mathrm{cross}}$. Panels 8, 8 ($\alpha=2,2/3$; $\theta/\pi=0.1$) show bulk-gap regime ($N_{\mathrm{cross}}=0$), while panels 8, 8 ($\alpha=2,2/3$; $\theta/\pi=0.4$) show edge-gap regime ($N_{\mathrm{cross}}\geq 1$). As with $\gamma_{N}$, the edge-bulk pattern remains qualitatively visible, though the phase diagram obtained by scanning the full parameter space $(\alpha,\theta)$ shows degraded sharpness compared to boundary seeds (compare Fig. 3 and Fig. 9). The recursion depth varies due to our stability criterion $b_{n}\lesssim 10^{-7}$, beyond which all coefficients are excluded (Appendix B.4).
@@
Figure 9: Phase diagrams for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries ($N=1000$) showing the crossing-count diagnostic $N_{\mathrm{cross}}(\alpha,\theta)$ (black: $N_{\mathrm{cross}}\geq 1$; white: $N_{\mathrm{cross}}=0$) for different seed operators: 9 $\gamma_{1}+\gamma_{2}$ (boundary seed), 9 $\gamma_{N}$ (bulk seed), 9 $\gamma_{N}+\gamma_{N+1}$ (bulk seed). Solid curves show the edge-bulk gap boundary $\Delta_{\mathrm{edge}}=\Delta_{\mathrm{bulk}}$ from the BdG spectrum for three edge-weight thresholds: $\omega_{\mathrm{edge}}=0.05,0.1,0.5$ (with $\ell_{\mathrm{edge}}=\lfloor\sqrt{N}\rfloor$). The boundary seed 9 shows the most quantitative agreement with the BdG-derived boundary (up to grid precision and finite-size), while bulk seeds 9, 9 couple to both edges and probe predominantly bulk-extended excitations, resulting in degraded quantitative matching even though qualitative edge-bulk trends remain visible. This demonstrates that boundary seeds are essential for obtaining a robust and quantitative Krylov diagnostic of edge-versus-bulk gap control. Data computed on a $99\times 99$ grid in $(\alpha,\theta)$ with $\alpha\in(0,3]$ and $\theta\in(0,\pi)$.

### v3

@@ C.4 Physical interpretation and topological criterion
Figure 7: BdG spectrum $E_{\nu}/\sqrt{\mathrm{Tr}(H^{2})}$ versus $\theta/\pi$ at extreme pairing-dominated $\epsilon=10$ for three long-range exponents $\alpha$ ($N=1000$, open boundaries). The systematic $\alpha$-dependence observed in Figs. 1 and 6 persists: gapless region for $\alpha=3$, similar gapless range for $\alpha=1$ with gap opening, and degeneracy lifted except for a small near-degenerate regime for $\alpha=1/3$. Spectral density near $E=0$ is further increased relative to smaller $\epsilon$, occurring uniformly across all $\alpha$ including the short-range limit. The qualitative $\alpha$-dependence remains unchanged across various $\epsilon$ regimes.
Appendix D Additional Results for Spectrum and the Role of Long-Range Pairing
@@
Figure 6 displays the spectrum at $\epsilon=1$. The qualitative $\alpha$-dependence observed at $\epsilon=-0.2$ (Fig. 1) persists: panel 6 ($\alpha=3$) exhibits a gapless region centered near $\theta/\pi\approx 0.5$ with sparse spectral density near $E=0$ elsewhere; panel 6 ($\alpha=1$) shows modified spectral density; and panel 6 ($\alpha=0.1$) displays the degeneracy lifted except for a small near-degenerate regime, with markedly denser spectral density near $E=0$ throughout. Compared to $\epsilon=-0.2$, the enhanced pairing amplitude at $\epsilon=1$ increases the density of modes near $E=0$ across all three $\alpha$ values.
Figure 7 presents the extreme pairing-dominated limit $\epsilon=10$. The systematic $\alpha$-dependence remains qualitatively unchanged: gapless region over a finite $\theta/\pi$ range for $\alpha=3$ (panel 7), similar gapless range with gap opening for $\alpha=1$ (panel 7), and degeneracy lifting to a small near-degenerate regime for $\alpha=1/3$ (panel 7). The spectral density near $E=0$ is further amplified relative to both $\epsilon=-0.2$ and $\epsilon=1$, reflecting the dominance of the pairing term $(1+\epsilon)\sin\theta$ in Eq. (6b). Notably, this increased density occurs uniformly across all $\alpha$ regimes, including the short-range limit where exponential localization would otherwise suppress accumulation of states near $E=0$.
@@
Figure 8: Lanczos coefficients $\{b_{n}\}$ for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries and boundary seed $\gamma_{1}+\gamma_{2}$ ($N=1000$). Each panel shows the two interleaved subsequences (odd and even recursion steps), whose relative ordering determines the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ and crossing count $N_{\mathrm{cross}}$. Panels 8 ($\alpha=2$, $\theta/\pi=0.1$) and 8 ($\alpha=2/3$, $\theta/\pi=0.1$) lie in the bulk-gap regime and exhibit no interchange of the two subsequences ($N_{\mathrm{cross}}=0$), while panels 8 ($\alpha=2$, $\theta/\pi=0.4$) and 8 ($\alpha=2/3$, $\theta/\pi=0.4$) lie in the edge-gap regime and show clear interchanges ($N_{\mathrm{cross}}\geq 1$). The qualitative pattern of no crossing in the bulk-gap phase as well as nonzero crossings in the edge-gap phase matches that observed for the single-operator seed $\gamma_{1}$ (Fig. 2), though the quantitative sharpness of the phase diagram when scanned over full parameter space $(\alpha,\theta)$ is reduced compared to the minimal boundary seed presented in the main text (compare Fig. 3 and Fig. 11). As in Fig. 2, the total number of Lanczos coefficients varies across parameter points. To ensure numerical stability, the recursion is terminated when $b_{n}\lesssim 10^{-7}$, and all subsequent coefficients are excluded from the analysis (see Appendix B.4 for details on stability checks).
Appendix E Alternative seed operators
@@
Figures 8, 9, and 10 show the Lanczos coefficients for these three seeds at representative parameter points spanning both short-range and long-range regimes, as well as bulk-gap and edge-gap phases. In all cases, the qualitative pattern observed in the main text persists: the two interleaved subsequences (odd and even recursion steps) do not interchange in the bulk-gap regime, yielding $N_{\mathrm{cross}}=0$, while clear interchanges occur in the edge-gap regime, producing $N_{\mathrm{cross}}\geq 1$. This confirms that the crossing-count signature is robust across seed choices and reflects genuine edge-bulk physics rather than an artifact of the specific seed operator.
However, when the crossing-count diagnostic is computed over the full parameter space $(\alpha,\theta)$, the quantitative sharpness of the phase boundary varies systematically with seed localization, as shown in Fig. 11. The boundary seed $\gamma_{1}+\gamma_{2}$ (panel 11) produces a phase diagram in close quantitative agreement with the BdG-derived edge-bulk gap boundary, though slightly less sharp than the minimal seed $\gamma_{1}$ used in the main text. By contrast, the bulk seeds $\gamma_{N}$ and $\gamma_{N}+\gamma_{N+1}$ (panels 11, 11) couple comparably to both chain ends and predominantly probe bulk-extended excitations, resulting in noticeably degraded quantitative matching with the BdG boundary even though the qualitative edge-bulk distinction remains visible. These results demonstrate that boundary-localized seeds are essential for obtaining a quantitatively reliable Krylov diagnostic of edge-versus-bulk gap control, while bulk seeds provide at best a qualitative probe.
@@
Figure 9: Lanczos coefficients $\{b_{n}\}$ for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries and bulk seed $\gamma_{N}$ ($N=1000$). Each panel shows the two interleaved subsequences (odd and even recursion steps), whose relative ordering determines the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ and crossing count $N_{\mathrm{cross}}$. Panels 9 ($\alpha=2$, $\theta/\pi=0.1$) and 9 ($\alpha=2/3$, $\theta/\pi=0.1$) lie in the bulk-gap regime ($N_{\mathrm{cross}}=0$), while panels 9 ($\alpha=2$, $\theta/\pi=0.4$) and 9 ($\alpha=2/3$, $\theta/\pi=0.4$) lie in the edge-gap regime ($N_{\mathrm{cross}}\geq 1$). The qualitative edge-bulk distinction persists for this bulk-localized seed, which couples comparably to both chain ends, though the phase diagram (obtained by scanning the full parameter space $(\alpha,\theta)$) sharpness degrades compared to boundary seeds (compare Fig. 3 and Fig. 11). To maintain numerical rigor, the recursion is terminated when $b_{n}\lesssim 10^{-7}$, excluding all subsequent coefficients (see Appendix B.4).
@@
Figure 10: Lanczos coefficients $\{b_{n}\}$ for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries and bulk seed $\gamma_{N}+\gamma_{N+1}$ ($N=1000$). Each panel shows the two interleaved subsequences whose relative ordering determines the staggering parameter $\eta_{n}=\ln(b_{2n-1}/b_{2n})$ and crossing count $N_{\mathrm{cross}}$. Panels 10, 10 ($\alpha=2,2/3$; $\theta/\pi=0.1$) show bulk-gap regime ($N_{\mathrm{cross}}=0$), while panels 10, 10 ($\alpha=2,2/3$; $\theta/\pi=0.4$) show edge-gap regime ($N_{\mathrm{cross}}\geq 1$). As with $\gamma_{N}$, the edge-bulk pattern remains qualitatively visible, though the phase diagram obtained by scanning the full parameter space $(\alpha,\theta)$ shows degraded sharpness compared to boundary seeds (compare Fig. 3 and Fig. 11). The recursion depth varies due to our stability criterion $b_{n}\lesssim 10^{-7}$, beyond which all coefficients are excluded (Appendix B.4).
@@
Figure 11: Phase diagrams for the long-range Kitaev chain at $\epsilon=-0.2$ with open boundaries ($N=1000$) showing the crossing-count diagnostic $N_{\mathrm{cross}}(\alpha,\theta)$ (black: $N_{\mathrm{cross}}\geq 1$; white: $N_{\mathrm{cross}}=0$) for different seed operators: 11 $\gamma_{1}+\gamma_{2}$ (boundary seed), 11 $\gamma_{N}$ (bulk seed), 11 $\gamma_{N}+\gamma_{N+1}$ (bulk seed). Solid curves show the edge-bulk gap boundary $\Delta_{\mathrm{edge}}=\Delta_{\mathrm{bulk}}$ from the BdG spectrum for three edge-weight thresholds: $\omega_{\mathrm{edge}}=0.05,0.1,0.5$ (with $\ell_{\mathrm{edge}}=\lfloor\sqrt{N}\rfloor$). The boundary seed 11 shows the most quantitative agreement with the BdG-derived boundary (up to grid precision and finite-size), while bulk seeds 11, 11 couple to both edges and probe predominantly bulk-extended excitations, resulting in degraded quantitative matching even though qualitative edge-bulk trends remain visible. This demonstrates that boundary seeds are essential for obtaining a robust and quantitative Krylov diagnostic of edge-versus-bulk gap control. Data computed on a $99\times 99$ grid in $(\alpha,\theta)$ with $\alpha\in(0,3]$ and $\theta\in(0,\pi)$.